UT 03 Jun 2021 - D/E first half, PEPSI 2nd half



Night Information

  • Sunset 19:29
  • Twilight Ends 21:03
  • Night Center 00:18
  • Handover to PEPSI 00:12 if using LUCI/MODS, handover 00:08 if using LBCs

Night Plan

  • Possible tasks:
    • TMS testing (half night)
      • Test 1:
        • Collimate and set reference
        • 1h in-focus data with TMS running in active mode
        • Collimate and note difference
      • Test 2:
        • Collimate (set new reference?)
        • Re-inject z4/z11:
          • LBCB z4 = -1050 and z11 = +600
          • LBCR z4 = -875 and z11 = +500
        • Take a series of extra-focal pupil images with active TMS.
    • MODS Field Correction Verification for IT 8315 - 1hr (POSTPONED)
      • check that collimation is stable
      • take 2 stars worth of field_collect, /IDL
    • LUCI-1 AO
      • LUCI-AO Suspended until further notice
      • On-axis ESM (Hoag's Object) - 1 hour
      • Crossover test (full AO w/N375 camera) 1-1.5 hours

Calibrations:

  • MODS Calibrations prior to opening and when closed for clouds. mods2[b|r].20210603.00NN.fits
cal m1b m1r m2b m2r
grlamps 1-3 1-3 1-3 1-3
grpixflats 4-8,9-13 4-8,9-13 4-8,9-13 4-8,9-13

Night Log:

19:30 Open

19:52 Slewing to a field for collimation

20:15 Starting collimation, dohybrid. It said it converged, but the last z4 was -1542 nm (and this was not just a delay in the monitor). The problem was that no good pupils were found in the image. The program should have backed out the Z4/Z11 corrections and exited, but instead it flagged the side with no pupils as "converged". This is a bug.

20:26 Repeating FPIA (started a while ago). One 1 or two pupils are found, although there are many in the image.

20:27 Trying again with /x2 to bring up the fainter pupils. So far, the estimated seeing is 2.2-2.8 and seeing is not being measured. 20:41 One more FPIA. Still not fitting z11.
  • Script to take in-focus images will be used and Z, Z4 and Z11 offsets made through PSF (as global offsets). No dithering and 99 30-sec exposures.
    • After FPIA, set Z global to have a -0.8 mm offset (global offests with this -0.8mm one are now -1.68)
    • set z4/z11 global offsets to -1050/600 on SX and -875/500 on DX
    • Slew to field and there set lbc_set_tms_offset.py -s 2
    • Now play OB.

21:00 - Taking the series of extra-focal pupil images.
  • Pupils are a tad too large, R=37 vs R=31. Need to adjust z global offset by +0.155 mm. (37.9 um of mirror Z motion corresponds to 1000 nm Z4).

21:17 or 21:18 will be the first images after the z global offset adjustments. (This discontinuity is easily seen in the FPIA output).

21:36 Stopped. Remove global offsets and run FPIA while I prepare script for in-focus asteroid observations.

21:57 Starting dofpia, /x2 on asteroid field. Finally the estimated seeing is low enough that Z11 and Z22 corrections are being made.

22:07 Slewing to field.
  • setting reference once first exposure is underway: set at 22:10
  • starting tms_loop.py at 22:10
  • on red side, 0.9" and on blue side about the same if not better (0.84")
  • 22:24 Clouds... images have no stars.
  • 22:32 stars are back - quick impression is that the red IQ has degraded but the blue is not so much worse as when we started.
  • Glass temp is still catching up to the ambient but closer: SX 11.2 vs 11.6 and on DX 11.05 vs 11.74. At the start of the night the difference was about 2 deg: 9.4 vs 11.6
  • 22:46 red 1.3" and blue 0.95"
  • 22:50 clouds coming back
23:03 Stop TMS and run dofpia, /x2
  • On DX, dz4 -2389 nm
  • On SX, dz4 555 nm
23:10 Starting long exposure on this field (3*39-sec)
  • take reference (lbc_set_tms_offset -s 2)
  • start active TMS loop
  • watching for how TMS handles z4 corrections from the focus chip and guide corrections. The latter should be accounted for already, the bigger question is about the former. First look - TMS seems to be backing out the tech chip focus corrections. (Now the question is how to allow z4 corrections to go through, if we do want them. And then, do we want them... do they make things better, worse or do nothing?)

2021/06/03 06:13:14.751220 S R TRACKERS            Active Optics Z04 +0.00000000 +0.00000000
2021/06/03 06:13:17.873399 S B TRACKERS            Active Optics Z04 +0.00000000 +0.00000000
2021/06/03 06:15:50.688640 S R TRACKERS            Active Optics Z04 +33.00000000 +0.00000000
2021/06/03 06:16:00.609354 S B TRACKERS            Active Optics Z04 +114.00000000 +0.00000000
2021/06/03 06:18:27.937288 S R TRACKERS            Active Optics Z04 -163.00000000 +0.00000000
2021/06/03 06:18:47.647526 S B TRACKERS            Active Optics Z04 +305.00000000 +0.00000000
2021/06/03 06:21:04.447449 S R TRACKERS            Active Optics Z04 +295.00000000 +0.00000000
2021/06/03 06:21:33.441829 S B TRACKERS            Active Optics Z04 +28.00000000 +0.00000000
2021/06/03 06:23:42.249249 S R TRACKERS            Active Optics Z04 -47.00000000 +0.00000000
2021/06/03 06:24:16.764342 S B TRACKERS            Active Optics Z04 +104.00000000 +0.00000000
2021/06/03 06:26:17.406954 S R TRACKERS            Active Optics Z04 +277.00000000 +0.00000000
2021/06/03 06:27:02.820923 S B TRACKERS            Active Optics Z04 +130.00000000 +0.00000000
2021/06/03 06:27:53.765019 S R TRACKERS            Active Optics Z04 -58.00000000 +0.00000000
2021/06/03 06:27:55.118529 S B TRACKERS            Active Optics Z04 +109.00000000 +0.00000000

23:28 Slewing to a new field, F3, while keeping the tms_loop running.
  • we lost the guide star on the blue side. One the red side it's a bit fuzzy
    • there was a jump on the first blue image 063156 but not on the first red one (063213)
    • got 2 pairs of exposures - 2nd blue on is better but also elongated.
    • John and Andrew looking at logs and discussing this jump --- is it related to the filtering? As we start the slew, the standard deviation is large - does it allow bad data through?
23:39 Copointing on this field. 23:47 sending copointing corrections (this is just for the sake of sending the copointing offsets). 23:49 Starting dofpia, /x2 on this field (F3). Stopped TMS shortly after I started dofpia.
  • Clouds - dofpia did not find pupils and exited
23:51 starting dofpia, /x2 again - we are looking at clouds.No pupils found. 23:57 restarted the tms_loop with the old reference and will let this go until handover. 00:02 dofpia, /x2
  • stopped tms soon after the loop was started.
  • There is a lot of spherical - both z11 and 22 - but FPIA took them out (wow! why didn't it do that the other night?)
  • FPIA converged just on time to handover to PEPSI, 00:08.

00:08 We can reconfigure to PEPSI+PFU

RB_Science crashed with this output. There were a lot of windows open on robs :

      object:focus field | filter:r-SLOAN | exptime:32.25 | pa_pnt:0.05 | dithoffx:0 | dithoffy:0
Pause? 1|2: 
----- receiving 'lbcb.20210603.033953.fits' [frame 1] (5.0s/105.3s)
      object:focus field | filter:V-BESSEL | exptime:32.28 | pa_pnt:0.06 | dithoffx:0 | dithoffy:0
Pause? 1|2: 
----- receiving 'lbcr.20210603.035430.fits' [frame 2] (5.0s/912.6s)
      object:F2 copoint st | filter:R-BESSEL | exptime:30.24 | pa_pnt:0.16 | dithoffx:0 | dithoffy:0
Pause? 1|2: 
----- receiving 'lbcb.20210603.035434.fits' [frame 1] (5.3s/907.7s)
      object:F2 copoint st | filter:V-BESSEL | exptime:30.27 | pa_pnt:0.16 | dithoffx:0 | dithoffy:0
Pause? 1|2: 
----- receiving 'lbcr.20210603.035532.fits' [frame 2] (5.0s/60.3s)
      object:F2 copoint st | filter:R-BESSEL | exptime:30.25 | pa_pnt:0.09 | dithoffx:0 | dithoffy:0
Pause? 1|2: 
----- receiving 'lbcb.20210603.035544.fits' [frame 1] (5.0s/70.1s)
      object:F2 copoint st | filter:V-BESSEL | exptime:30.29 | pa_pnt:0.11 | dithoffx:0 | dithoffy:0
Pause? 1|2: 
----- receiving 'lbcr.20210603.035635.fits' [frame 2] (5.0s/61.3s)
      object:F2 copoint st | filter:R-BESSEL | exptime:30.24 | pa_pnt:0.10 | dithoffx:0 | dithoffy:0
Pause? 1|2: 
----- receiving 'lbcb.20210603.035654.fits' [frame 1] (5.0s/70.4s)
      object:F2 copoint st | filter:V-BESSEL | exptime:30.27 | pa_pnt:0.11 | dithoffx:0 | dithoffy:0
Pause? 1|2: 
----- receiving 'lbcr.20210603.035737.fits' [frame 2] (5.0s/65.5s)
      object:F2 copoint st | filter:R-BESSEL | exptime:30.25 | pa_pnt:0.10 | dithoffx:0 | dithoffy:0

Unknown error! Please restart (and report the following messages).
------------------------------------------------------------
Traceback (most recent call last):
  File "/lbt/lbto/supportscripts/ObserverSupport/RB_Science.py", line 951, in display
    sched.run()
  File "/lbt/lbto/supportscripts/ObserverSupport/RB_Science.py", line 611, in run
    t.run()
  File "/lbt/lbto/supportscripts/ObserverSupport/RB_Science.py", line 860, in run
    [f'frame {self.__frameno}']+self.__cmds.expand(img))
  File "/lbt/lbto/supportscripts/ObserverSupport/RB_Science.py", line 353, in set
    raise err # temporarily rethrow the exception for now.
  File "/lbt/lbto/supportscripts/ObserverSupport/RB_Science.py", line 351, in set
    self.__ds9.set(cmd)
  File "/usr/lib64/python3.7/site-packages/pyds9/pyds9.py", line 658, in set
    s, blen, 1)
  File "/usr/lib64/python3.7/site-packages/pyds9/xpa.py", line 170, in xpaset
    raise ValueError(errmsg)
ValueError: XPA$ERROR: no response from server during handshake (?:?)

------------------------------------------------------------
^C^CException ignored in: <module 'threading' from '/usr/lib64/python3.7/threading.py'>
Traceback (most recent call last):
  File "/usr/lib64/python3.7/threading.py", line 1307, in _shutdown
    lock.acquire()
  File "/lbt/lbto/supportscripts/ObserverSupport/RB_Science.py", line 55, in exit_handler
    sys.exit(0)
SystemExit: 0


-- BarryRothberg - 28 May 2021
Topic revision: r11 - 03 Jun 2021, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback